# Fast LiDAR Informed Visual Search in Unseen Indoor Environments

### v8

@@ III-D Map Updates
Two global occupancy maps (search maps) are maintained. One for the LiDAR and the other for the visual sensor. Global search maps are implemented using ROS Costmap2D [costmap]. All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$. The robot is equipped with a visual sensor with a cone shaped FoV, implemented as a triangular costmap. The LiDAR sensor has a 360 degree FoV, represented as a square-type costmap. Observations for both sensors are a set of range estimates to the nearest occupied point $e\in\mathcal{E}^{\text{occ}}$. Points between the robot and the nearest occupied point along each ray of the LiDAR scan are free points $e\in\mathcal{E}^{\text{free}}$. Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$. This process is shown in Fig. 2(d).
III-E Multisensor Frontier Exploration
The planning module seeks to generate a viewpoint at each planning step that results in the highest likelihood to find the target of interest. This is achieved in several steps: candidate viewpoints are generated, viewpoints are scored, and finally the point with the highest score is fed to the robot for navigation to that point.
@@ III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [yamauchi1997frontier], shown in Fig. 2(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Any group of frontier points beyond a minimum size becomes its own frontier. Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering. Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.
@@ III-E2 Viewpoint Selection
Candidate viewpoints are selected based on a computationally efficient heuristic utility function. High value viewpoints are those that are near the robot, allow inspection of many unknown map points and result in inspection of non-map points. This process is shown in Fig. 2(f). The utility functions is a weighted sum of the following: (1) Penalize distance from robot position to viewpoint, (2) Reward expected number of unknown cells discovered, (3) Reward viewpoints with frontiers near the path to the candidate viewpoint, and (4) Reward viewpoints that result in inspection of a scaled number non-map points. Scaling parameters on each term are tuned during experiments. Weight are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered. The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
@@ IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on a real robot in unseen indoor environments.
Map Target NBVP RRT Multisensor Frontier Ground Truth LIVES
Name Difficulty [bircher2016receding] [rrt] Exploration [vidal2020multisensor] Classification (Our Method)
@@
Table I: Average time (s) for each method to find the object of interest in the given simulation environments.
@@ IV-B Simulation and Baseline Comparison
Simulations are performed in Gazebo with the Turtlebot [tbot]. Move base is used to navigate to the given viewpoints [navigation-ros]. Baselines are: 1) a modified Multisensor Frontier Exploration [vidal2020multisensor], 2) the Next-Best View Planner (NBVP) [bircher2016receding], 3) a Rapidly-Exploring Random Tree (RRT) [rrt], and 4) LIVES given ground-truth scan classifications. We make the following modifications to ref. [vidal2020multisensor] for our environments. The method is projected into 2D and viewpoints are sampled at centroids instead of surrounding them. This modified multisensor frontier method is equivalent to the proposed approach without segmented scan information. Two Gazebo worlds are tested with easy and hard setups in each. A figure depicting them is shown on the project website. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES. The easy setting has fewer objects in the environment while hard has more objects. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position. Easy and hard settings are defined by the search object in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. In the hard setting LIVES and Ground-Truth perform nearly identically and show a 15% improvement with the addition of non-map points compared with the Multisensor method. There exists a large gap in the easy office setting with a near 30% improvement using contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. The results indicate that LIVES outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
@@ IV-C Hardware Experiments
Experiments are in a 20mx30m apartment and a 15mx38m hallway environment. These two environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot is equipped with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera. The search object is one or two small rolling suitcases. YOLOv5 [yolov5] is used for detection. Objects are localized by first converting PointCloud data into the optical frame and then converting to pixel coordinates using camera intrinsics. Bounding boxes are then used to obtain the depth information from coinciding pixels. A time lapse of a simple trials in the apartment is shown in Fig. 5. The search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the candidate on the left hand side.
Figure 5: A time lapse of a simple scenario. The ground truth map (20x30m) is shown in the lower right corner (a). This is unavailable to the robot and not included in the training dataset. In (a), the global search map, frontier points and non-map points are also shown. The target object is behind the robot starting position. The result of the first planning step is also shown. The planner takes advantage of the concentration of non-map points behind the robot. Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
Figure 6: A time lapse of the hallway environment (15x38m) experiment scenario. In the lower right corner of (a) is the ground truth map with the suitcase locations and non-map points near the robot shown. This information is not in the training dataset and not given to the robot. In (a) the robot pose is a centered on the visual costmap. The first pose selected inspects the non-map points behind the robot. In (b) the robot chooses to turn left at the hallway split, recognizing a trash can as a non-map point. This correspondence can be seen between the non-map point in (b) appearing in the viewpoint visited by the robot in (c). In (d) the robot selects the circled viewpoint, facing to the left. This results in viewpoint (e), inspecting the suitcase, which was recognized as a non-map points in (d). In (f) the robot has just localized the first suitcase and takes the decision to return to the fork for non-map points found by that candidate viewpoint, leading to the second suitcase. (g) shows the overall robot path overlaid on the ground truth map.
@@
However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby, resulting in detection. We compare the accuracy of the trained classifier in these out of distribution environments to ground truth information. The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home. The video demonstrates three trials. The first and second take place in the apartment with one and two search objects, respectively. In both apartment scenarios, the robot tends to look towards furniture like couches, shelves and tables. Similarly, Spot also tends to inspect trash cans and other objects in the hallway. This behavior demonstrates the impact of non-map points and suggests the efficacy of the segmentation policy. Several key moments in the hallway search scenario are shown in the time lapse in Fig. 6.
@@ V Conclusions
We introduce LIVES, a method to leverage contextual information for search in unseen environments. It is shown to outperform baselines by 10-30% in simulation. Through ablative studies, we validate our architecture and training approach for map-free scan classification. Our experiments confirm the effectiveness of the visual search planner and learning method in real-world scenarios. Ongoing work plans to expand to multi-robot applications.

### v11

@@ III-D Map Updates
Two global occupancy maps (search maps) are maintained. One for the LiDAR and the other for the visual sensor (see fig. 3(a-b)). Global search maps are implemented as occupancy grids. They are updated at each step from local sensor occupancy grids, which are implemented using the ROS Costmap2D package [costmap]. All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$. The robot is equipped with a visual sensor with a cone shaped FoV, implemented as a triangular costmap. The LiDAR sensor has a 360 degree FoV, represented as a square-type costmap. Observations for both sensors are a set of range estimates to the nearest occupied point $e\in\mathcal{E}^{\text{occ}}$. Points between the robot and the nearest occupied point along each ray of the LiDAR scan are free points $e\in\mathcal{E}^{\text{free}}$. Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$.
III-E Viewpoint Planning
The planner sends the viewpoint at each planning step that results in the highest likelihood to find the target of interest. This involves picking viewpoints that balance exploration of unknown space in the environment with exploiting non-map points from LiDAR scans. This is achieved in several steps: candidate viewpoints are generated, viewpoints are scored, and best sample is fed to the robot.
@@ III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [yamauchi1997frontier], shown in Fig. 3(c-d). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Any group of frontier points beyond a minimum size becomes its own frontier. Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering. Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor (see fig. 4).
Figure 4: An overview of the viewpoint sampling process. Four viewpoints are considered at each centroid, shown as red triangles. The highest scoring viewpoint is shown as a yellow triangle to be sent to the robot.
@@ III-E2 Viewpoint Selection
Candidate viewpoints are scored using the following utility function. High value viewpoints are those that are near the robot, allow inspection of many unknown map points and result in inspection of non-map points. The utility functions is a weighted sum of the following: (1) Penalize distance from robot position to viewpoint, (2) Reward expected number of unknown cells discovered, (3) Reward viewpoints with frontiers near the path to the candidate viewpoint, and (4) Reward viewpoints that result in inspection of non-map points. Scaling parameters on each term are tuned during experiments. Weights are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered. The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
@@ IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) it reduces detection time compared against existing algorithms, and (3) it can be deployed successfully on a real robot in unseen indoor environments.
Map Target NBVP RRT MFE Ground Truth Our
Name Difficulty [bircher2016receding] [umari2017autonomous] [vidal2020multisensor] Classification Method
@@
Table I: Average time in seconds (and success rate) for each method to find the object of interest in simulation environments.
@@ IV-B Simulation and Baseline Comparison
Simulations are performed in Gazebo with the Turtlebot and ROS navigation. Baselines are: 1) modified Multisensor Frontier Exploration (MFE) as described in section II [vidal2020multisensor], 2) Next-Best View Planner (NBVP) [bircher2016receding], 3) Rapidly-Exploring Random Tree (RRT) [umari2017autonomous], and 4) our method, but informed with ground-truth scan classifications as described in section III-B.
Figure 6: Environments used for baseline comparison. Robot start pose is denoted by a cyan circle. (a) The Apartment (20x30m) simulation environment in the hard configuration. Easy configuration has fewer objects. (b) The Office (25x45m) environment. Easy setup involves the search target to the left of the red line and hard involves the search target on the right.
@@
Two Gazebo worlds are tested with easy and hard setups in each (fig. 6). The apartment (20x30m) examines the impact of non-target objects on the planner, the easy setting with fewer objects. The office (25x45m) tests the method in a large environment under significantly varied target position. Easy and hard settings are defined by the search object in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
Overall, the results support the proposed approach, reporting 100% success and outperforming baselines across all evaluation settings. Further, the nearly identical performance between the proposed approach with both predicted and ground truth labels support the efficacy of the perception module. While MFE performed similarly to the proposed approach in the easy apartment, the lack of the labeled scan data slowed task completion by 18-42% across the remaining settings. NBVP and RRT baselines struggled in most settings, with the former averaging about 50% success. RRT outperformed other baselines in the hard office environment, however, was 20% less successful and moderately slower than the proposed approach. The overall results support our method, suggesting the benefits of focusing on non-map points. This focus allows the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
Figure 7: A toy experimental scenario. (a) The first planning step. Candidate viewpoints are sampled at yellow disks. The robot chooses the next viewpoint at the shown cyan arrow in order to visually inspect the concentration of non-map points behind the robot’s starting pose. The robot pose on the ground truth map is shown in the lower right (unknown to the robot, not included in the training dataset). (b) The concentration of non-map points is found to be two red chairs and the search object (suitcase). This result reinforces that the scan classifier successfully recognizes non-permanent features in unseen indoor environments.
@@ IV-C Hardware Experiments
Experiments are in a 20mx30m apartment and a 15mx38m hallway environment. These two environments are not included in dataset $\mathcal{D}$ used to train the classification policy. A Boston Dynamics Spot is equipped with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera. We first evaluate the accuracy of the trained classifier in these out of distribution environments to ground truth information. This is accomplished by teleoperating the robot while running the ground truth and map-free classifiers in parallel. The accuracy in the two out of distribution environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These suggests the classifier’s ability to distinguish permanent features common to many indoor environments.
During experiments, YOLOv5 [yolov5] is used for detection. Objects are localized by first converting PointCloud data into the optical frame and then converting to pixel coordinates using camera intrinsics. Bounding boxes are then used to obtain the depth information from coinciding pixels. A simple experiment is shown in Fig. 7 and demonstrates the benefits of non-map information in guiding the vision sensor in the robot’s surroundings.
Figure 8: Paths the robot took to find the objects of interest (suitcases) in the video scenarios 2 and 3. (a) Two search objects in the 20x30m apartment environment. (b) Two search objects in the 15x38m hallway environment.
@@
Instead of selecting based on number of unknown cells, contextual information leads the agent to inspect many non-map features nearby, resulting in detection. Videos of experimentation, implementation details, code, training data, and all maps can be found on the project website https://sites.google.com/view/lives-2024/home. The video shows two experiments in the apartment and one in the hallway. Fig. 8 demonstrates the path of the robot in scenarios 2 and 3. In both apartment scenarios, the robot tends to look towards furniture like couches, shelves and tables. Similarly, Spot also tends to inspect trash cans and other objects in the hallway. This behavior demonstrates the impact of non-map points and further supports the efficacy of the segmentation policy. In the hallway environment the robot is guided by non-map points at several key forks. After first inspecting the concentration of non-map points situated behind the robot, the decisions the robot makes at the two key forks it intersects lead it to quickly find the two suitcases in the environment. Overall, the results of hardware experiments on the Spot support hypothesis (3).
@@ V Conclusions
We introduce a novel visual planner that leverages contextual information found from wide field-of-view LiDAR scans. The presented map-free LiDAR classifier identifies non-permanent points in the environment for the planner’s attention. The method is shown to outperform baselines by 10-30% in simulation. Through ablative studies, we validate our architecture and training approach for map-free scan classification. Our experiments confirm the effectiveness of the visual search planner and learning method in real-world scenarios. Future work plans to leverage the non-map points to generate higher quality maps and expand to multi-robot applications.